# Bayesian Signal Matching for Transfer Learning in ERP-Based Brain Computer Interface

### v2

@@ 1.2 Challenges and Existing Work
Transfer learning based on Riemannian geometry has gained increasing attention recently due to its fast speed to converge and a natural framework to leverage information from source participants. The input data for RG were the sample covariance matrices, and the distance-based algorithm based on RG was called Minimum Distance to Mean (MDM) (Barachant et al.,, 2011). Congedo et al., and Barachant and Congedo, extended the original RG-based classifier to the P300 ERP design by augmenting the covariance matrix with cross-covariance between observed signals and reference target signals. Rodrigues et al., applied the Riemannian procrustes analysis (RPA) to address the heterogeneity of EEG signals across different sessions or participants. Before the authors applied the MDM classifier, they applied certain affine transformations to raw participant-level covariance matrices such that the resulting covariance matrices were less heterogeneous across sessions or participants while their Riemannian distances were preserved. Li et al., standardized the covariance matrices across participants by applying the affine transformation with the participant-specific Riemannian geometric mean covariance matrix. Finally, Khazem et al., proposed Minimum Distance to Weighted Mean (MDWM) by combining estimated mean covariance matrices from source participants and the new participant via the Riemannian distance with a trade-off hyper-parameter.
@@ 2.1 Basic Notations and Problem Setup
Let $n=0,\ldots,N$ be the participant index, where $n=0$ refers to the new participant and $n=1,\ldots,N$ refer to source participants. Let $l=1,\ldots,L_{n}$ and $i=1,\ldots,I_{n}$ be the character index and sequence index for participant $n$, respectively. We follow the conventional RCP design such that each sequence contains $J(J=12)$ stimuli, including six row stimuli from top to bottom $(1,\ldots,6)$ and six column stimuli from left to right $(7,\ldots,12)$ on the $6\times 6$ virtual keyboard (See the virtual screen in Figure 1). For the $i$th sequence, $l$th target character, and $n$th participant, let $\bm{W}_{n,l,i}=(W_{n,l,i,1},\ldots,W_{n,l,i,12})^{\top}$ be a stimulus code indicator that takes values from the permutation of $\{1,\ldots,12\}$. Under the RCP design, given the target character and the stimulus-code indicators, there are exactly two target stimuli and ten non-target stimuli within each sequence. We define $\bm{Y}_{n,l,i}=(Y_{n,l,i,1},\ldots,Y_{n,l,i,12})^{\top}$ as the stimulus-type indicator for the $i$th sequence, $l$th target character, and $n$th participant, where $Y_{n,l,i,j}\in\{0,1\}$. For example, given a target character “T” and a stimulus code indicator $\bm{W}_{n,l,i}=(7,9,10,5,1,2,8,11,6,4,3,12)^{\top}$, we obtain its stimulus type indicator $\bm{Y}_{n,l,i}=(0,0,0,0,0,0,1,0,0,1,0,0)^{\top}$, where the indices of $1$s correspond to the indices of the $4$th row and $2$nd column (denoted by $8$) in $\bm{W}_{n,l,i}$. Let $k=0,\ldots,K-1$ be the cluster index, where $k=0$ is the cluster that matches the new participant and $k=1,\ldots,K-1$ are the clusters within source participants. Finally, we incorporate $E$ channels of EEG signals and let $e(e=1,\ldots,E)$ be the channel index. We extract channel-specific EEG segments from the onset of each stimulus with a long EEG response window length (with $T_{0}$ time points). We denote $\bm{X}_{n,l,i,j,e}(s)$ as the extracted EEG signal segment of the $j$th stimulus, $i$th sequence, $l$th target character, and $n$th participant from channel $e$ at time $s\in[0,s_{T_{0}}]$.
@@ 2.2 The Framework of Bayesian Signal Matching
We made three assumptions to our proposed framework as follows: (i). We assume that P300 ERPs are the same regardless of target character to spell. To simplify, we let all source participants spell the same character $\omega$ with the same sequence replication size $I$, so we drop the target character index $l$ and the source participant index $n$ from $I_{n}$. The biological mechanism behind the speller system is the response to an unexpected stimulus. Under the RCP design, there were always two target stimuli supposed to elicit target P300 ERPs, and this rule applied regardless of the spelling characters. (ii). We assumed that target and non-target ERP functions shared the same covariance matrix within the same cluster. We simplified the temporal covariance matrix as a scaled AR(1) correlation matrix. Linear discriminant analysis (LDA) with its variations have been shown to be effective binary classifiers for this problem, we argued that the difference on the mean level had sufficient separation effect, and sharing the same covariance matrix reduced the number of parameters. Empirical correlation matrix also suggested that the correlation declined fast and they were similar for target and non-target stimuli, even after the xDAWN filter. The shared and simplified covariance matrix also reduced the number of parameters and made the proposed model parsimonious, especially when we considered borrowing data from source participants. (iii). We performed selection with respect to the target ERP data and we excluded non-target ERP data from the source participants’ pool. Despite a larger sample size of non-target ERP data, unfortunately, their signals showed fewer unique patterns and would add more noises to the model fitting. The right panel of Figure 1 provides a detailed illustration of the data integration framework of proposed Bayesian signal matching method. Details of three assumptions can be found in the supplementary material Section X.
@@ 3.1 Model Reparametrization and Prior Specifications
$\operatorname{vec}\left\{\left[\bm{f}(s_{1}),\cdots,\bm{f}(s_{T_{0}})\right]\right\}\sim\mathcal{MVN}\left(\operatorname{vec}\left(\bm{M}\right),\bm{\Sigma}\otimes\bm{\Omega}\right),$ (6)
where $\bm{f},\bm{\mu}\in\mathbb{R}^{E}$ are column vectors whose components are the functions $\{\bm{f}_{e}\}_{e=1}^{E}$ and $\{\bm{\mu}_{e}\}_{e=1}^{E}$, respectively. Furthermore, $\bm{M}\in\mathbb{R}^{E\times T_{0}}$ with $\bm{M}_{e,t}=\bm{\mu}_{e}(s_{t})$, $\bm{\Sigma}\in\mathbb{R}^{T_{0}\times T_{0}}$ with $\bm{\Sigma}_{t_{1},t_{2}}=k(s_{t_{1}},s_{t_{2}})$, where $t_{1},t_{2}\in\mathbb{Z}$ and $0\leq t_{1},t_{2}\leq T_{0}$, and $\otimes$ is the Kronecker product operator. We denote $\bm{f}\sim\mathcal{MGP}(\bm{\mu},\kappa,\bm{\Omega})$.
In our case, we let $\bm{b}_{n,1}$ be an $E$-dim target ERP function of participant $n$ characterized by $\bm{\mu}_{n,1},\kappa_{1},\bm{\Sigma}_{n}^{s}$, $n=0,\cdots,N$, and $\bm{b}_{0,0}$ be another $E$-dim non-target ERP function of participant 00 characterized by $\bm{\mu}_{0,0},\kappa_{0},\bm{\Sigma}_{0}^{s}$. To simplify, we let $\bm{\mu}_{n,1}=\bm{\mu}_{0,0}=\bm{0}$ and apply a $\gamma$-exponential kernel function to specify target kernel $\kappa_{1}$ and non-target kernel $\kappa_{0}$ as follows:
$k(z_{i},z_{j})=\psi_{0}\exp\left\{-\left(\frac{||z_{i}-z_{j}||_{2}^{2}}{s_{0}}\right)^{\gamma_{0}}\right\},$ (7)
@@
$\psi_{n,1},\psi_{0,0}\overset{\mathrm{indep.}}{\sim}\mathcal{LN}(0,1),\quad\sigma_{n,e}\sim\mathcal{HC}(0,5.0),\quad\rho_{n}\sim\mathcal{U}(0,1),\quad\eta_{n}\sim\mathcal{U}(0,1),$
@@ 3.4 Posterior Character-Level Prediction
[…]
@@ 4.1 Naive Multi-channel Simulation Scenario
Figure 3: The left and right panels show the means and standard errors of testing prediction accuracy of Cases 1 and 2 by BSM-Mixture, BSM, BSM-Reference, MDWM, and swLDA across 100 replications, respectively. The upper, middle, and lower panels show prediction accuracy with the first 2, 3, and 4 training sequence replications, respectively.
@@ 4.2 Real Data Based Simulation Scenario
ResultsFor matching, we only observe that participant 1 has a posterior mean of being merged around 0.2, while the posterior means of other source participants are below 0.1. The finding is consistent with the matching from the real data analysis given the same participant ordering and supports our argument that we typically expect one or two matching based on BSM classifier. Details of selection probability can be found in Section S4.2 in the supplementary. For prediction, Figure 4 shows the testing prediction accuracy by the same five methods across 100 replications. The upper, middle, and lower panel are the prediction accuracy by training the first 2, 3, and 4 training sequences of the new participant. Overall, we observe that BSM-Mixture performs better than MDWM and swLDA. The prediction curves of BSM-Mixture and BSM are almost identical, suggesting that the algorithm identifies at least one source participant with the posterior mean of being merged exceeds the threshold $\delta_{Z}=0.1$, which is also consistent with the matching result. The prediction accuracy of BSM is slightly higher and narrower than that of BSM-Reference, possibly because the borrowing information may be limited when only one participant has reached the selection threshold. Finally, when we increase the sample size of training data, the prediction accuracy improves accordingly.
@@ 6 Discussion
For future work, first, we would replace the generative base classifier with other robust discriminant classifiers, e.g., Riemannian geometry. The entire framework would remain unchanged except that we would specify the likelihood function by introducing Riemannian Gaussian distributions (Zanini et al.,, 2016), (Said et al.,, 2017). Even if we followed the generative pathway for classifiers, despite the poor performance of the original SMGP, it could be further improved by a slight modification. Since the original SMGP method focused on inferring on spatial-temporal brain activity under the P300 ERP-BCI setting rather than prediction, it usually required a large sample size for model fitting. Therefore, if we continued using the current generative Bayesian classifier, we would adopt the idea of split-and-merge to explicitly control the discrepancies between target and non-target (transformed) EEG signals, i.e., the temporal interval close to 800 ms on the stimulus level. Finally, we would incorporate our framework into calibration-free methods (Verhoeven et al.,, 2017). In addition to borrowing the training data from source participants and testing separately, we would directly start testing on the new participant and dynamically update the parameters of the baseline group as the data arrive.

### v3

@@ 1.2 Challenges and Existing Work
Transfer learning based on Riemannian geometry (RG) has gained increasing attention recently due to its fast speed to converge and a natural framework to leverage information from source participants. The input data for RG were the sample covariance matrices, and the distance-based algorithm based on RG was called Minimum Distance to Mean (MDM) (Barachant et al.,, 2011). Congedo et al., and Barachant and Congedo, extended the original RG-based classifier to the P300 ERP design by augmenting the covariance matrix with cross-covariance between observed signals and reference target signals. Rodrigues et al., applied the Riemannian procrustes analysis (RPA) to address the heterogeneity of EEG signals across different sessions or participants. Before the authors applied the MDM classifier, they applied certain affine transformations to raw participant-level covariance matrices such that the resulting covariance matrices were less heterogeneous across sessions or participants while their Riemannian distances were preserved. Li et al., standardized the covariance matrices across participants by applying the affine transformation with the participant-specific Riemannian geometric mean covariance matrix. Finally, Khazem et al., proposed Minimum Distance to Weighted Mean (MDWM) by combining estimated mean covariance matrices from source participants and the new participant via the Riemannian distance with a trade-off hyper-parameter.
@@ 2.1 Basic Notations and Problem Setup
Let $n=0,\ldots,N$ be the participant index, where $n=0$ refers to the new participant and $n=1,\ldots,N$ refer to source participants. Let $l=1,\ldots,L_{n}$ and $i=1,\ldots,I_{n}$ be the character index and sequence index for participant $n$, respectively. We follow the conventional RCP design such that each sequence contains $J(J=12)$ stimuli, including six row stimuli from top to bottom $(1,\ldots,6)$ and six column stimuli from left to right $(7,\ldots,12)$ on the $6\times 6$ virtual keyboard (See the virtual screen in Figure 1). For the $i$th sequence, $l$th target character, and $n$th participant, let $\bm{W}_{n,l,i}=(W_{n,l,i,1},\ldots,W_{n,l,i,12})^{\top}$ be a stimulus code indicator that takes values from the permutation of $\{1,\ldots,12\}$. Under the RCP design, given the target character and the stimulus-code indicators, there are exactly two target stimuli and ten non-target stimuli within each sequence. We define $\bm{Y}_{n,l,i}=(Y_{n,l,i,1},\ldots,Y_{n,l,i,12})^{\top}$ as the stimulus-type indicator for the $i$th sequence, $l$th target character, and $n$th participant, where $Y_{n,l,i,j}\in\{0,1\}$. For example, given a target character “T” and a stimulus code indicator $\bm{W}_{n,l,i}=(7,9,10,5,1,2,8,11,6,4,3,12)^{\top}$, we obtain its stimulus type indicator $\bm{Y}_{n,l,i}=(0,0,0,0,0,0,1,0,0,1,0,0)^{\top}$, where the indices of $1$s correspond to the indices of the $4$th row and 22nd column (denoted by $8$) in $\bm{W}_{n,l,i}$. Let $k=0,\ldots,K-1$ be the cluster index, where $k=0$ is the cluster that matches the new participant and $k=1,\ldots,K-1$ are the clusters within source participants. Finally, we incorporate $E$ channels of EEG signals and let $e(e=1,\ldots,E)$ be the channel index. We extract channel-specific EEG segments from the onset of each stimulus with a long EEG response window length (with $T_{0}$ time points). We denote $\bm{X}_{n,l,i,j,e}(s)$ as the extracted EEG signal segment of the $j$th stimulus, $i$th sequence, $l$th target character, and $n$th participant from channel $e$ at time $s\in[0,s_{T_{0}}]$.
@@ 2.2 The Framework of Bayesian Signal Matching
We make three assumptions as follows: (i). We assume that P300 ERPs are the same regardless of target character to spell. To simplify, we let all source participants spell the same character $\omega$ with the same sequence replication size $I$, so we drop the target character index $l$ and the source participant index $n$ from $I_{n}$. For example, $Y_{n,l,i,j}$ is reduced to $Y_{n,i,j}$. The biological mechanism behind the speller system is the response to an unexpected stimulus. Under the RCP design, there were always two target stimuli supposed to elicit target P300 ERPs, and this rule applied regardless of the spelling characters. (ii). We assumed that target and non-target ERP functions shared the same covariance matrix within the same cluster. We simplified the temporal covariance matrix as a scaled AR(1) correlation matrix. Linear discriminant analysis (LDA) with its variations have been shown to be effective binary classifiers for this problem, we argued that the difference on the mean level had sufficient separation effect, and sharing the same covariance matrix reduced the number of parameters. Empirical correlation matrix also suggested that the correlation declined fast and they were similar for target and non-target stimuli, even after the xDAWN filter. The shared and simplified covariance matrix also reduced the number of parameters and made the proposed model parsimonious, especially when we considered borrowing data from source participants. (iii). We performed selection with respect to the target ERP data and we excluded non-target ERP data from the source participants’ pool. Despite a larger sample size of non-target ERP data, unfortunately, their signals showed fewer unique patterns and would add more noises to the model fitting. The right panel of Figure 1 provides a detailed illustration of the data integration framework of proposed Bayesian signal matching method. Details of three assumptions can be found in the supplementary material Section S1.
@@ 3.1 Model Reparametrization and Prior Specifications
$\operatorname{vec}\left\{\left[\hat{\bm{f}}(s_{1}),\cdots,\hat{\bm{f}}(s_{T_{0}})\right]\right\}\sim\mathcal{MVN}\left(\operatorname{vec}\left\{\left[\hat{\bm{\mu}}(s_{1}),\cdots,\hat{\bm{\mu}}(s_{T_{0}})\right]\right\},\bm{\Sigma}\otimes\bm{\Omega}\right),$ (6)
where $\operatorname{vec}$ was the vectorization operator; $\hat{\bm{f}}(s),\hat{\bm{\mu}}(s)\in\mathbb{R}^{E}$ were column vectors that evaluated the functions $\bm{f}$ and $\bm{\mu}$ at the index point $s$. Furthermore, $\bm{\Sigma}\in\mathbb{R}^{T_{0}\times T_{0}}$ with $\bm{\Sigma}_{t_{1},t_{2}}=k(s_{t_{1}},s_{t_{2}})$, where $t_{1},t_{2}\in\mathbb{Z}$ and $0<t_{1},t_{2}\leq T_{0}$, and $\otimes$ is the Kronecker product operator. We denote $\bm{f}\sim\mathcal{MGP}(\bm{\mu},\kappa,\bm{\Omega})$.
In our case, we let $\bm{b}_{n,1}$ be an $E$-dim target ERP function of participant $n$ characterized by $\bm{\mu}_{n,1},\kappa_{1},\bm{\Sigma}_{n}^{s}$, $n=0,\cdots,N$, and $\bm{b}_{0,0}$ be another $E$-dim non-target ERP function of participant 0 characterized by $\bm{\mu}_{0,0},\kappa_{0},\bm{\Sigma}_{0}^{s}$. To simplify, we let $\bm{\mu}_{n,1}=\bm{\mu}_{0,0}=\bm{0}$ and apply a $\gamma$-exponential kernel function to specify target kernel $\kappa_{1}$ and non-target kernel $\kappa_{0}$ as follows:
@@
where $0\leq\gamma_{0}<2,s_{0}>0$. In practice, $\gamma_{0}$ and $s_{0}$ are treated as hyper-parameters and selected among a group of values and we do not further differentiate among participants. For participant-specific kernel variance parameters $\psi_{n,1}$ and $\psi_{0,0}$, we assign Log-Normal priors with the a mean zero and a scale one. In addition, we define a function $\phi(\cdot)$ as an eigenfunction of kernel $k$ with eigenvalue $\lambda$ if it satisfies the following integral: $\int k(s,s^{\prime})\phi(s)ds=\lambda\phi(s^{\prime}).$ In general, the number of eigenfunctions is infinite, i.e., $\{\phi_{l}(s)\}_{l=1}^{+\infty}$, and we rank them by the descending order of their corresponding eigenvalues $\{\lambda_{l}\}_{l=1}^{+\infty}$. The Mercer’s Theorem suggests that the kernel function $k$ can be decomposed as a weighted summation of normalized eigenfunctions $\phi_{l}(s)$ with positive eigenvalues $\lambda_{l}$ and approximated with a sufficient large $L$ subgroup as follows: $k(s,s^{\prime})=\sum_{l=1}^{+\infty}\lambda_{l}\phi_{l}(s)\phi_{l}(s^{\prime})\approx\sum_{l=1}^{L}\lambda_{l}\phi_{l}(s)\phi_{l}(s^{\prime})$. Thus, a random function $\bm{f}$ can be represented as a linear combination of the above eigenfunctions and approximated with a sufficient large $L$ subgroup evaluated at $s$ as follows:
@@
[…]
@@ 3.4 Posterior Character-Level Prediction
$\operatorname{Pr}(\bm{Y}_{0}^{*}=\bm{y}^{*}_{\omega}\mid\bm{X}_{0}^{*},\bm{W}_{0}^{*},\bm{X},\bm{W},\bm{Y})$ $=\int\operatorname{Pr}(\bm{Y}_{0}^{*}=\bm{y}^{*}_{\omega}\mid\bm{\Theta}_{0};\bm{X}_{0}^{*},\bm{W}_{0}^{*})\operatorname{Pr}(\bm{\Theta}_{0}\mid\bm{X},\bm{W},\bm{Y})d\bm{\Theta}_{0},$ (11)
@@ 4.1 Naive Multi-channel Simulation Scenario
Figure 3: The left and right panels show the means and standard errors of testing prediction accuracy of Cases 1 (Unmatched) and 2 (Matched) by BSM-Mixture, BSM, BSM-Reference, MDWM, and swLDA across 100 replications, respectively. The upper, middle, and lower panels show prediction accuracy with the first 2, 3, and 4 training sequence replications, respectively. For Unmatched scenarios, BSM-Mixture was robust to source participants by adopting BSM-Reference’s results, while for Matched scenario, BSM-Mixture required one testing sequence fewer than BSM-Reference to achieve the 90% threshold with narrower error bars.
@@ 4.2 Real Data Based Simulation Scenario
Results  For matching, we only observe that participant 1 has a posterior mean of being merged around 0.2, while the posterior means of other source participants are below 0.1. The finding is consistent with the matching from the real data analysis given the same participant ordering and supports our argument that we typically expect one or two matching based on BSM classifier. Details of selection probability can be found in Section S4.2 in the supplementary. For prediction, Figure 4 shows the testing prediction accuracy by the same five methods across 100 replications. The upper, middle, and lower panels are the prediction accuracy by training the first 2, 3, and 4 training sequences of the new participant. Overall, we observe that BSM-Mixture performs better than MDWM and swLDA. The prediction curves of BSM-Mixture and BSM are almost identical, suggesting that the algorithm identifies at least one source participant with the posterior mean of being merged exceeds the threshold $\delta_{Z}=0.1$, which is also consistent with the matching result. The prediction accuracy of BSM is slightly higher and narrower than that of BSM-Reference, possibly because the borrowing information may be limited when only one participant has reached the selection threshold. Finally, when we increase the sample size of training data, the prediction accuracy improves accordingly.
@@ 6 Discussion
For future work, first, we would replace the generative base classifier with other robust discriminant classifiers, e.g., Riemannian geometry. The entire framework would remain unchanged except that we would specify the likelihood function by introducing Riemannian Gaussian distributions (Zanini et al.,, 2016), (Said et al.,, 2017). If we followed the generative pathway for classifiers, it could be further improved byadopting the split-and-merge concept to explicitly control the discrepancies between target and non-target (transformed) EEG signals. Finally, we would incorporate our framework into calibration-free methods (Verhoeven et al.,, 2017). In addition to borrowing the training data from source participants and testing separately, we would directly start testing on the new participant and dynamically update the parameters of the baseline group as the data arrive.
Data Available Statement Since the participants in the original study did not consent to release their data in public directly, even for de-identified version. However, the de-identified data are available upon request for other researchers. An outgoing data user agreement (DUA) will be required to initiate the data transfer. Please contact the corresponding author for details.
7 Acknowledgments
The authors would like to thank the Editor Professor Michael Stein, the Associate Editor and reviewers for their helpful comments and constructive suggestions, which led to this much-improved manuscript. The authors would also like to acknowledge the participants in our BCI experiments. The content is solely the responsibility of the authors and does not represent the official viewpoints of NICHD, NIH, NIDRR, or the Department of Education.
8 Disclosures
This work was partially supported by grants NIH R01DA048993 (Kang and Johnson), NIH R01MH105561 (Guo and Kang), and NSF IIS2123777 (Kang). The authors have no conflicts of interests to declare.